# Sheet Music Transformer: End-To-End Optical Music Recognition Beyond Monophonic Transcription

### v1

@@ 1 Introduction
The progress made in OMR, which initially depended on multi-stage workflows [27], has diversified with the emergence of Deep Learning solutions. Two major approaches currently exist: object detection workflows [36, 5],—in which notes are individually detected and assembled into a digital document—and holistic workflows [32, 8, 4, 34, 2], aka end-to-end methods, in which systems directly generate a symbolic representation of a region of a given document. The latter paradigm dominates the current state of the art of other applications, such as the recognition of text, speech or mathematical formulae [11, 10, 38], and has also proven successful in OMR.
@@
The remainder of the paper is structured as follows: in Section 2, we present the polyphonic transcription challenge in OMR, explain the current state of the art and discuss its limitations. In Section 3, we describe the end-to-end SMT Neural Network proposed in this paper, while the experiments conducted are presented in Section 4, along with the results and analysis that can be drawn from them. Finally, we conclude the paper in Section 7, in addition to discussing other avenues for future work.
@@ 3.2 Decoder
[…]
@@ 3.3 Output Encoding
Figure 2: Example of an excperpt from Humdrum **kern-encoded pianoform music. The blue dashed line represents a **kern spine, which is a voice in the musical score, in this case, a staff. The green box represents simultaneous notes during interpretation, which are represented as a text line in the ground truth. The red box is a single musical symbol (in this case, a note). The ground truth is read top to bottom and left to right, while the score is read from left to right and bottom to top.
@@ 4 Experimental setup
In this section, we present the experimental framework designed in order to evaluate the performance of our method in comparison to other state-of-the–art methods [^5] Public immplementation of the experiments is given in https://github.com/multiscore/SMT.
@@ 4.1.1 GrandStaff
The GrandStaff dataset is a publicly available corpus [^6] https://sites.google.com/view/multiscore-project/datasets [29] that consists of $53,882$ printed images of single-line (or system) pianoform scores, along with their digital score encoding. The dataset is composed of both original works from six authors—from the Humdrum [^7] https://github.com/humdrum-tools/humdrum-data repository—and synthetic augmentations of the music encodings that make it possible to provide a greater variety of musical sequences and patterns. The dataset comes with an official partition, in which the $7,000$ original scores are used as a test set, and the $46,882$ samples generated from augmentations comprise the training set. The dataset comes with an alternative version that introduces distortions into images in order to make them resemble low quality photocopies. This version is, from here on, referred to as Camera GrandStaff. Figure 3 depicts two samples of this dataset.

### v2

@@ 1 Introduction
The progress made in OMR, which initially depended on multi-stage workflows [27], has diversified with the emergence of Deep Learning solutions. Two major approaches currently exist: object detection workflows [36, 5],—in which notes are individually detected and assembled into a digital document—and holistic workflows [32, 8, 4, 34, 2], also known as end-to-end methods, in which systems directly generate a symbolic representation of a region of a given document. The latter paradigm dominates the current state of the art of other applications, such as the recognition of text, speech or mathematical formulae [11, 10, 38], and has also proven successful in OMR.
@@
The remainder of the paper is structured as follows: in Section 2, we present the polyphonic transcription challenge in OMR, explain the current state of the art and discuss its limitations. In Section 3, we describe the end-to-end SMT Neural Network proposed in this paper, while the experiments conducted are presented in Section 4, along with the results, Section 5, and analysis, Section 6, that are drawn from them. Finally, we conclude the paper in Section 7, in addition to discussing other avenues for future work.
@@ 3.2 Decoder
The decoder consists of a language model that, at each time step $t$ (which is conditioned by the input feature vector produced by the encoder, $x_{e}^{\prime}$ along with the previously predicted tokens, $\left(\hat{y}_{0},\ldots,\hat{y}_{t-1}\right)$) outputs a probability distribution $p_{t}\in\mathbb{R}^{\left|\Sigma\right|}$ over the $\Sigma$ vocabulary symbols, with the predicted token $\hat{y}_{t}$ being that which maximizes this probability score. The process starts with a special start-of-transcription symbol, $>\hat{y}_{0}=\text{\textless{sot}\textgreater}$, and continues until an end-of-transcription token is predicted, $>\hat{y}_{|\mathbf{\hat{y}}|}=\text{\textless{eot}\textgreater}$.
@@ 3.3 Output Encoding
Figure 2: Example of an excerpt from Humdrum **kern-encoded pianoform music. The blue dashed line represents a **kern spine, which is a voice in the musical score, in this case, a staff. The green box represents simultaneous notes during interpretation, which are represented as a text line in the ground truth. The red box is a single musical symbol (in this case, a note). The ground truth is read top to bottom and left to right, while the score is read from left to right and bottom to top.
@@ 4 Experimental setup
In this section, we present the experimental framework designed in order to evaluate the performance of our method in comparison to other state-of-the-art methods [^5] The implementation of the model and the datasets are available in https://grfia.dlsi.ua.es/sheet-music-transformer/.
@@ 4.1.1 GrandStaff
The GrandStaff dataset is a publicly available corpus [^6] https://sites.google.com/view/multiscore-project/datasets [29] that consists of $53,882$ printed images of single-line (or system) pianoform scores, along with their digital score encoding. The dataset is composed of both original works from six authors—from the Humdrum [^7] https://github.com/humdrum-tools/humdrum-data repository—and synthetic augmentations of the music encodings that make it possible to provide a greater variety of musical sequences and patterns. The dataset comes with an official partition, in which the $7,000$ original scores are used as a test set, and the $46,882$ samples generated from augmentations comprise the training set. The dataset comes with an alternative version that introduces distortions into images to make them resemble low-quality photocopies. This version is, from here on, referred to as Camera GrandStaff. Figure 3 depicts two samples of this dataset.
@@ 5 Results
6 Discussion
Although the results described in the previous section are positive, an analysis of the SMT performance is required in order to explain and understand these improvements. This study is carried using the model that performed best in our experiments: the $\text{SMT}_{\text{NexT}}$.
One of the most noteworthy results is the substantial improvement as regards the LER metric. For example, the $\text{SMT}_{\text{CNN}}$, despite producing slightly worse CER and SER than the CRNN baseline, notably improves the LER. Although this metric suggests an overall improvement to the output quality, it is difficult to assess what this implies. In order to clarify this, we evaluated what percentage of the test results had a correct **kern structure. That is, we assessed how many documents could be processed by a standard musicological tool, such as Verovio Humdrum Viewer [26] or MuseScore. The results of this method are depicted in Table 2.
Table 2: Average Render percentage (%) values for the best-performing models for both the baseline and our approach in the test sets for the collections evaluated.
Samples CRNN $\text{SMT}_{\text{NexT}}$
LER Render % LER Render %
GrandStaff 7661 23.2 70.3 13.1 98.6
Camera GrandStaff 7661 29.5 51.1 13.5 96.4
Quartets 6107 67.0 22.5 5.6 90.7
These results make it possible to state that there is a clear correlation between LER and overall document quality, as an improvement to this metric significantly increases the number of directly usable documents, which can be directly edited and processed with musicological tools. It is consequently possible to state that the SMT produces outputs that are easier for end-users to edit and handle, thus making them more usable.
Two visualization errors from both the Camera GrandStaff and Quartets dataset are shown in Figures 5 and 7, respectively. The highlighted errors were computed using the musicdiff tool [16] and visualized with the MuseScore editor [^11] https://musescore.org/es.
(a) Input test image
(b) Render of the prediction made by the $\text{SMT}_{\text{NexT}}$ with highlighted transcription errors.
Figure 5: Test example from the GrandStaff dataset with the errors highlighted. This specific sample attained a CER of 5.0%, a SER of 6.0% and a LER of 20.5%.
Both examples show that, visually, the model is capable of generating a correct music sequence in terms of syntax, as no bar-completion or time-related errors are found. Indeed, most of the errors that are highlighted in these examples show that most of them are pitch misplacements. This means that, although the model is able to recognize the shape of the notes accurately, it generally fails to predict their position in their staff, or to place accidentals. Another interesting case is that shown in Figure 6.
Figure 6: Example of a redundant error in the Camera GrandStaff dataset, in which the ground truth contains redundant annotations that the language model avoids.
[…]
(a) Input test image
(b) Prediction from the $\text{SMT}_{\text{NexT}}$.
Figure 7: Test example from the Quartets dataset with the errors highlighted. This specific sample attained a CER of 1.0%, a SER of 1.8% and a LER of 3.2%.
This is a specific case in music notation in which text-based metrics penalize unfairly, as the piece produced is musically correct and equivalent to the ground truth. It is, therefore, possible that the metrics presented in Table 5, although necessary to assess document-quality information, are pessimistic figures of merit for the methodological evaluation. Even though no consensus has been reached as to the best way in which to evaluate OMR, and that this is still, therefore, an open research question, we corroborate that the results of this paper are robust, usable and musically-accurate, and this is also supported by the information provided in Table 2.
7 Conclusion
In this paper, we tackle the challenge of advancing Optical Music Recognition beyond monophonic transcription, which has traditionally involved simplifications or ad-hoc adaptations of existing methods. We introduce the Sheet Music Transformer (SMT) model, an autoregressive Transformer-based model that utilizes attention computation and language modeling to transcribe input scores into digital music encodings. We evaluate this approach in two polyphonic music scenarios: pianoform and string quartet scores. Our results demonstrate that the SMT model not only effectively transcribes complex musical layouts, but also outperforms current state-of-the-art methods, thus implying a significant advance in OMR.
This research also opens avenues for further improvements to OMR. Firstly, and as discussed in Section 6, standard OMR evaluation methods may overlook musicological interpretations, leading to pessimistic results. Addressing this issue could involve exploring graphic-based output music encodings [7] or employing additional musicological metrics in order to assess transcription accuracy [25]. The development of segmentation-free full-page transcription methods is another promising direction, as it is not limited by image features or layout constraints. Finally, we propose that the Universal OMR transcription challenge, a desired goal in OMR research, could be addressed by profiting from the language modeling capabilities of the SMT, thus allowing it to be trained to transcribe music engraved in various manners.
8 Acknowledgements
This paper is part of the project I+D+i PID2020-118447RA-I00 (MultiScore), funded by MCIN/AEI/10.13039/501100011033. The first author is supported by grants ACIF/2021/356 and CIBEFP/2022/19 from the “Programa I+D+i de la Generalitat Valenciana”.